OpenStudy (anonymous):

Simplify (4/3y-2)-(5/9y^2-4)

OpenStudy (campbell_st):

need a common denominator or (3y-2)(3y+2)\[\frac{4(3y+2)}{(3y-2)(3y+2)} - \frac{5}{(3y-2)(3y+2)} = \frac{12y +3}{9y^2 - 4}\]

OpenStudy (anonymous):